my daughter asked how old i am, i answered, 'in 20 years i'll be twice as old as i was 12 years ago.'How old am i now?
Answer by Edwin McCravy(20086)


my daughter asked how old i am, i answered, 'in 20 years i'll be twice
as old as i was 12 years ago.'How old am i now?

1. First I'll tell you the answer.  
2. Then I'll show you how to check it.
3. Then I'll tell you how to find it.


1. The answer is "I am 44 now"

2. I am 44 years old.  In 20 years I'll be 64. 12 years ago I was 32.  
   And 64 = 2 times 32.

3. I am x years old.  In 20 years I'll be x+20.  12 years ago I was x-12.
   And x+20 = 2 times x-12.

   So:

   x + 20 = 2(x - 12)

Solve that and you'll get 44.

Do you see that from 2 and 3 that often setting up a word problem
is exactly the same process as checking it when you know the answer?
Just "check" it using x.  Then solve.

Edwin